please help me solve this equation 2log x - log 5 = log 125
Answer by saifoo.khan(162)

2log x - log 5 = log 125
log x^2 - log 5 = log 125
lg( x^2/5) = lg 125
x^2/5 = 125
x^2 = 125*5
x^2 = 626
taking square root on both sides,

x = sqrt(625)
x = 25
answer. =)

-----------------------
If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($150/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($1.5-$3.0 per problem). Contact me at einstein.saifoo@gmail.com
Thank you. =)